OpenStudy (anonymous):

Translate the sentence into an inequality. Six times the sum of a number and 28 is at least−29. Use the variable y for the unknown number. Will medal

OpenStudy (kittiwitti1):

Six times: x6 Sum of number y and 28: y+28\[\text{Is at least} -29\rightarrow\ge-29\]Put everything together and you get\[6\times(y+28)\ge-29\]

OpenStudy (kittiwitti1):

Or:\[6(28+y)\ge-29\]

OpenStudy (anonymous):

both are inccorect

OpenStudy (kittiwitti1):

hmmm.

OpenStudy (kittiwitti1):

at least is... OH Flip the sign.\[6(28+y)\le-29\]